OpenStudy (anonymous):

what is the simplified form of x times the square root of the quantity 64 x cubed,plus 2 x times the square root of x,plus 8 times the square root of the quantity 2 x

OpenStudy (anonymous):

\[x \sqrt{64} x^{3} + 2x \sqrt{x} + 8\sqrt{2x}\] you can simplify root of 64 to 8 since 8 times 8 = 64. for the other roots write them as \[x^{1/2}\] use properties of exponents to combine the x terms

OpenStudy (anonymous):

??